# Weighted Envy Freeness With Limited Subsidies

### v1

@@ 3.2 Computing a WEF-able allocation
For $t\in[T]$, Algorithm 1 computes a maximum-value matching $\mu^{t}$ in $H[I,O_{t}]$ and allocates the items $\mu_{i_{1}}^{t}\ldots\mu_{i_{w_{i}}}^{t}$ to agent $i$. By maximality of the matching, the items assigned in round $t$ are at least as valuable as the items remaining after that round, that is, $v_{i}(o_{1})\geq v_{i}(o_{2})$ for each $i\in N$, $o_{1}\in A_{i}^{t}$ and $o_{2}\in O_{t+1}$. Thus, for each round $t\in[T]$, for $i\neq j\in N$ and $o\in O_{t+1}$ that $i$ values the most, it holds that
$\frac{v_{i}(A_{i}^{t})}{w_{i}}=\frac{v_{i}(\{\mu_{i_{1}}^{t}\cup\ldots\cup\mu_{i_{w_{i}}}^{t}\})}{w_{i}}\geq\frac{w_{i}v_{i}(o)}{w_{i}}=$
$\frac{w_{j}v_{i}(o)}{w_{j}}\geq\frac{v_{i}(\{\mu_{j_{1}}^{t+1}\cup\ldots\cup\mu_{j_{w_{j}}}^{t+1}\})}{w_{j}}=\frac{v_{i}(A_{j}^{t+1})}{w_{j}}$
Therefore,
$\frac{v_{i}(A_{i})}{w_{i}}=\sum_{t=1}^{T}\frac{v_{i}(A_{i}^{t})}{w_{i}}\geq\sum_{t=1}^{T-1}\frac{v_{i}(A_{i}^{t})}{w_{i}}\geq\sum_{t=2}^{T}\frac{v_{i}(A_{j}^{t})}{w_{j}}=\frac{v_{i}(A_{j})}{w_{j}}-\frac{v_{i}(A_{j}^{1})}{w_{j}}$
Hence, the envy of agent $i$ for agent $j$ is bounded by
$\frac{v_{i}(A_{j})}{w_{j}}-\frac{v_{i}(A_{i})}{w_{i}}\leq\frac{v_{i}(A_{j}^{1})}{w_{j}}\leq\frac{w_{j}V}{w_{j}}=V.\qed$
Proposition 3.11 implies an upper bound of $(n-1)(W-w_{1})V$ on the total subsidy, but we can show an even better bound by bounding the subsidy per agent. First we generalize Lemma 4.1 of Brustle et al. [7] to the weighted setting:
Lemma 3.12.
Let $A$ be a WEF-able allocation. For any positive number $v$, If $cost_{A}(i,k)\geq-v$ for every arc $(i,k)$ in $G_{A,w}$, then the maximum subsidy required is at most $w_{i}v$ per agent $i\in N$.
Proof.
Assume $\ell_{i}(A)=(i\ldots j)$.
@@
$cost_{A}(C)=cost_{A}(\ell_{i}(A))+cost_{A}(j,i).$
By Theorem 3.2, $cost_{A}(C)\leq 0$, thus,
$cost_{A}(\ell_{i}(A))\leq-cost_{A}(j,i)\leq V.$
Therefore, $s_{i}=w_{i}cost_{A}(\ell_{i}(A))\leq w_{i}V$. ∎
In most allocations, including the one resulting from Algorithm 1, it is not always true that $cost_{A}(i,k)\geq-V$ for every arc $(i,k)$. We apply the technique from Brustle et al. [7] to use Lemma 3.12 with a modified valuation function $\bar{v}_{i}$, derived from the weighted valuation $\frac{v_{i}(A_{i})}{w_{i}}$.
Let $A^{t}$ be the output allocation from Algorithm 1, computed in iteration $t$. For each $i\in N$, we define the modified valuation function as follow:
@@
Under the modified valuations, for any two agents $i,j\in N$, the modified-cost assigned to the arc $(i,j)$ in the envy graph (with unit weights) is defined as $\overline{cost}_{A}(i,j)=\bar{v}_{i}(A_{i})-\bar{v}_{i}(A_{j})$. Moreover, the modified-cost of a path $(i_{1},...,i_{k})$ is $\overline{cost}_{A}(i_{1},...,i_{k})=\sum_{j=1}^{k-1}\overline{cost}_{A}(i_{j},i_{j+1})$.
We will prove that an allocation that is WEF-able for the original valuations is also WEF-able for the modified valuations (Proposition 3.14), and that the maximum subsidy required by each agent for the original valuations is bounded by the subsidy required for the modified valuations (Proposition 3.15).
Next, we demonstrate that under the modified valuations, the cost of each arc is at least $-V$. Finally, by Lemma 3.12, we conclude that the maximum subsidy required for any agent $i\in N$ is $w_{i}V$ for the modified valuation $\bar{v}$ (3.16) and for the original valuations $v$ as well.
Observation 3.13.
@@
By Theorem 3.2, it is sufficient to prove that all directed cycles in the envy graph (with the modified valuations and unit weights) have non-positive total cost. We prove a stronger claim: in every round $t$, the total modified-cost added to every directed cycle $C$ is non-positive. Let $A^{t}$ be the allocation computed by Algorithm 1 at iteration $t$. Suppose, contrary to our assumption, that there exists a cycle $C=(i_{1},...,i_{r})$ and a round $t$ in which the modified-cost added to $C$ is positive. To simplify notation, we consider $i_{1}$ as $i_{r+1}$. This implies that
@@
Case 1: All arcs $i_{j}\to i_{j+1}$ in $C$ have $\bar{v}_{i_{j}}(A_{i_{j+1}}^{t})=\frac{v_{i_{j}}(A_{i_{j+1}}^{t})}{w_{i_{j+1}}}$ (in particular, this holds for $t=T$). In this case, inequality (1) implies
$\overline{cost}_{A^{t}}(C)=\sum_{j=1}^{r}\frac{v_{i_{j}}(A_{i_{j}}^{t})}{w_{i_{j+1}}}-\frac{v_{i_{j}}(A_{i_{j}}^{t})}{w_{i_{j}}}>0$
Combined with Theorem 3.2, this contradicts Proposition 3.10, which states that $A^{t}$ is WEF-able.
Case 2: All arcs $i_{j}\to i_{j+1}$ in $C$ have $\bar{v}_{i_{j}}(A_{i_{j+1}}^{t})=\frac{v_{i_{j}}(A_{i_{j}}^{t+1})}{w_{i_{j}}}$. In this case, inequality (1) implies $\sum_{j=1}^{r}\frac{v_{i_{j}}(A_{i_{j}}^{t+1})}{w_{i_{j}}}>\sum_{j=1}^{r}\frac{v_{i_{j}}(A_{i_{j}}^{t})}{w_{i_{j}}}$. Notice that all the items in $A_{j_{1}}^{t+1},...,A_{j_{r}}^{t+1}$ are available at iteration $t$, which contradicts the optimality of $\mu^{t}$.
Case 3: Some arcs $i_{j}\to i_{j+1}$ in $C$ satisfy Case 1 and the other arcs satisfy Case 2. Let $l\geq 1$ be the number of arcs in $C$ that satisfy Case 2. We decompose $C$ into a sequence of $l$ edge-disjoint paths, denoted $P_{1},...,P_{l}$, such that the last node of each path is the first node of the next path, and in each path, only the last arc satisfies Case 2. Formally, suppose that some path contains $k\geq 1$ agents, denoted as $i_{1},...,i_{k}$, and $k-1$ arcs. Then for each $21\leq j\leq k-$ 2, $\bar{v}_{i_{j}}\left(A_{i_{j+1}}^{t}\right)=\frac{v_{i_{j}}\left(A_{i_{j+1}}^{t}\right)}{w_{i_{j+1}}}$ and $\bar{v}_{i_{k-1}}\left(A_{i_{k}}^{t}\right)=\frac{v_{i_{k-1}}\left(A_{i_{k-1}}^{t+1}\right)}{w_{i_{k-1}}}$. Since $\overline{cost}_{A^{t}}(C)>0$, there exits a path $P=(i_{1},...,i_{k})$ where $\overline{cost}_{A^{t}}(P)>0$, which implies that:
$0<\sum_{j=1}^{k-1}\left(\bar{v}_{i_{j}}(A_{i_{j+1}}^{t})-\bar{v}_{i_{j}}(A_{i_{j}}^{t})\right)=$
$\sum_{j=1}^{k-2}\left(\frac{v_{i_{j}}(A_{i_{j+1}}^{t})}{w_{i_{j+1}}}-\frac{v_{i_{j}}(A_{i_{j}}^{t})}{w_{i_{j}}}\right)+\frac{v_{i_{k-1}}(A_{i_{k-1}}^{t+1})}{w_{i_{k-1}}}-\frac{v_{i_{k-1}}(A_{i_{k-1}}^{t})}{w_{i_{k-1}}}$
The rest of the proof is similar to the proof of Proposition 3.10. We construct another allocation $B^{t}$ randomly as follows:
@@
$\mathbb{E}\left[\sum_{i\in N}\left(v_{i}(B_{i}^{t})-v_{i}(A^{t}_{i})\right)\right]=$
$\sum_{1\leq j\leq k-2}\left(\mathbb{E}\left[v_{i_{j}}(o_{i_{j+1}}^{t})\right]-\mathbb{E}\left[v_{i_{j}}(o_{i_{j}}^{t})\right]\right)+$
$\mathbb{E}\left[v_{i_{k-1}}(o_{i_{k-1}}^{t+1})\right]-\mathbb{E}\left[v_{i_{k-1}}(o_{i_{k-1}}^{t})\right]=$
$\sum_{1\leq j\leq k-2}\frac{v_{i_{j}}(A^{t}_{i_{j+1}})}{w_{i_{j+1}}}-\frac{v_{i_{j}}(A^{t}_{i_{j}})}{w_{i_{j}}}+\frac{v_{i_{k-1}}(A_{i_{k-1}}^{t+1})}{w_{i_{k-1}}}-\frac{v_{i_{k-1}}(A_{i_{k-1}}^{t})}{w_{i_{k-1}}}$
@@
Given Observation 3.13, for each $i,j\in N$,
$\bar{v}_{i}(A_{j})-\bar{v}_{i}(A_{i})\geq\frac{v_{i}(A_{j})}{w_{j}}-\frac{v_{i}(A_{i})}{w_{i}}$
@@
By Proposition 3.14, the allocation $A$ is WEF-able under the valuations $\bar{v}$. Together with Lemma 3.12, if for each $i,j\in N$ it holds that $\bar{v}_{i}(A_{j})-\bar{v}_{i}(A_{i})\geq-V$, the subsidy required for agent $i\in N$ is at most $w_{i}V$ for $\bar{v}$.
@@
$\sum_{t\in[T-1]}\max\left(\frac{v_{i}(A_{j}^{t})}{w_{j}},\frac{v_{i}(A_{i}^{t+1})}{w_{i}}\right)+\frac{v_{i}(A_{j}^{T})}{w_{j}}-\sum_{t\in[T]}\frac{v_{i}(A_{i}^{t})}{w_{i}}\geq$
@@
We are now prepared to prove the main theorem.
Theorem 3.17.
For additive valuations and integer entitlements, Algorithm 1 computes in polynomial time a WEF-able allocation where the subsidy to each agent is at most $w_{i}V$ and the total subsidy is at most $(W-w_{1})V$.
@@
Algorithm 1 clearly runs in polynomial-time in $n$ 222it is polynomial in $W$, as the graph $H$ contains $W$ copies. We can compute the max-value matching using reduction to network flow, in time polynomial in $n$..
From Proposition 3.10, we know that $A$ is WEF-able under the original valuations. Combined with Proposition 3.14 and Proposition 3.16 , $A$ is also WEF-able under the modified valuations and requires a subsidy of at most $w_{i}V$ for each agent $i\in N$.
By Proposition 3.15, it follows that under the original valuations, the required subsidy for each agent $i\in N$ is at most $w_{i}V$. From part (4) of Theorem 3.4, there is at least one agent who requires no subsidy, so the required total subsidy is at most $(W-w_{1})V$. ∎
@@
Brustle et al. [7] proved that the allocation resulting from their algorithm satisfies both EF and EF1. However, in the weighted setup, a WEF-able allocation may not satisfy $WEF(x,y)$ for any $x+y=1$. This also holds for the allocation produced by Algorithm 1 too.
Proposition 3.18.
For general additive valuations, a WEF-able allocation does not necessarily satisfy $WEF(x,y)$ for any $x+y=1$.
Proof.
Suppose 2 agents and 3 items with weights $w_{1}=1,w_{1}=10$ and valuations
$\begin{bmatrix}&1&2&3\\ 1&1&1&1\\ 2&10&10&10\end{bmatrix}$
The allocation $A_{1}=\emptyset,A_{2}=\{o_{1},o_{2},o_{3}\}$ is WEF-able since the cost of the cycle between the two agents in the weighted envy graph is $\frac{3}{10}-0+0-\frac{30}{10}<0$. However, if agent 1 receives at least one item, the cost of the cycle between the two agents in the weighted envy graph becomes at least $\frac{2}{10}-1+10-\frac{20}{10}>0$.
Thus, the only WEF-able allocation is $A_{1}=\emptyset,A_{2}=\{o_{1},o_{2},o_{3}\}$, but this violates WEF$(x,y)$ w.r.t. agent 2. ∎
@@ 4 WEF Solutions for Additive Identical Valuations
In this section we assume that all agents have the same additive valuation functions, denoted by $v$.
@@
The value per unit entitlement of each agent $i<n$ is $V(w_{i}-1)/w_{i}$, and for agent $n$ it is $V$. Therefore, to avoid envy, each agent $i<n$ should receive a subsidy of
$w_{i}\left(1-\frac{w_{i}-1}{w_{i}}\right)V=V$
and the total subsidy required is
$(n-1)V.$
@@
Observation 4.2.
Given an instance with identical valuations, let $A$ be any allocation, and denote by $(i,\ldots,j)$ any path in the envy-graph of $A$ between agents $i,j\in N$. Then,
$cost_{A}(i,\ldots,j)=\sum_{k=i}^{j-1}\left(\frac{v(A_{k+1})}{w_{k+1}}-\frac{v(A_{k})}{w_{k}}\right)=\frac{v(A_{j})}{w_{j}}-\frac{v(A_{i})}{w_{i}}.$
Using Observation 4.2, it is easy to prove the following lemma.
Lemma 4.3.
With identical valuations, every allocation is WEF-able.
@@
$\begin{split}&cost_{A}(C)=\\ &\sum_{k=1}^{r-1}\left(\frac{v(A_{i_{k+1}})}{w_{i_{k+1}}}-\frac{v(A_{i_{k}})}{w_{i_{k}}}\right)+\frac{v(A_{i_{1}})}{w_{i_{1}}}-\frac{v(A_{i_{r}})}{w_{i_{r}}}=\\ &\sum_{k=1}^{r}\left(\frac{v(A_{i_{k}}}{w_{i_{k}}}-\frac{v(A_{i_{k}}}{w_{i_{k}}}\right)=0\end{split}$ (2)
Hence, by condition (b) of Theorem 3.2, $A$ is WEF-able. ∎
@@
It first orders the items in $M$ based on their common value, such that $v(o_{1})\geq v(o_{2})\geq...\geq v(o_{m})$.
The algorithm then iteratively traverses the items. At iteration $t$ it selects the agent that minimizes the expression $\frac{v(A_{i}^{t-1}\cup{\{o_{t}\}})}{w_{i}}$, with ties are broken in favor of the agent with the higher index, which corresponds to the agent with the larger $w_{i}$.
Intuitively, this selection minimizes the likelihood that weighted envy is generated. Subsequently, the algorithm allocates item $o_{t}$ to the selected agent.
Example 4.4.
@@
    In order to mitigate envy, $s_{i_{1}}=\frac{6}{7}$ and $s_{i_{2}}=0$.
1:Instance $(N,M,v,\{w_{i}\}_{i\in N})$ with additive identical valuations.
2:WEF-able allocation $A=(A_{i})_{i\in N}$ with total required subsidy of at most $(n-1)V$.
3:Sort the items based on their value: $v(o_{1})\geq v(o_{2})\geq...\geq v(o_{m})$
4:$A_{i}\leftarrow\emptyset,\forall i\in N$
5:for $t=1$ to $m$ do
6:     $I^{t}\leftarrow\operatorname*{arg\,min}_{i\in N}\left(\frac{v(A_{i}^{t-1}\cup\{o_{t}\})}{w_{i}}\right)$
7:     $i^{t}\leftarrow\max_{i\in I^{t}}\left(i\right)$
8:     $A_{i^{t}}^{t}\leftarrow A_{i^{t}}^{t-1}\cup\{o_{t}\}$
9:     $A_{j}^{t}\leftarrow A_{j}^{t-1}$, for all $j\in N\textbackslash\{i^{t}\}$
10:end for
11:return $A$
Example 4.4 illustrates that the resulting allocation may not meet the WEF(1,0) criterion – Algorithm 2 might allocate all items to the agent with the highest entitlement. However, it does satisfy the WEF(0,1) criterion. The following proposition holds based on the selection rule, which ensures that the cost for any agent $i\in N$ towards the agent who received the new item is bounded by $\frac{V}{w_{i}}$.
Proposition 4.5.
@@
We prove by induction that at each iteration $t\in[T]$, where $o_{t}$ is the item assigned in this iteration and $i^{t}$ is the agent receiving this item, $A^{t}$ satisfies $WEF(0,1)$. The claim is straightforward for the first iteration. We assume the claim holds for the $(t-1)$-th iteration and prove it for the $t$-th iteration. Every agent $j\neq i^{t}$ is not envied due to the induction step. For $j=i^{t}$, due to the selection rule, $\frac{v(A^{t}_{i})+v(o^{t})}{w_{i}}=\frac{v(A^{t-1}_{i})+v(o^{t})}{w_{i}}\geq\frac{v(A^{t-1}_{j})+v(o^{t})}{w_{j}}=\frac{v(A^{t}_{j})}{w_{j}}$. This is exactly the definition of WEF$(0,1)$ for the pair $i,j$. ∎
Based on Observation 4.2 and Proposition 4.5, we show that each path starting at agent $i\in N$ under $WEF(0,1)$ allocation is bounded by $\frac{V}{w_{i}}$.
Proposition 4.6.
With identical additive valuations, for every WEF$(0,1)$ allocation $A$, $cost_{A}(\ell_{i}(A))\leq\frac{V}{w_{i}}$, for all $i\in N$.
@@
For each agent $i\in N$, denote the most costly path starting at $i$ in that graph by $\ell_{i}(A)=(i,...,j)$ for some agent $j\in N$. Then by Observation 4.2, $cost_{A}(\ell_{i}(A))=\frac{v(A_{j})}{w_{j}}-\frac{v(A_{i})}{w_{i}}$. By definition of WEF$(0,1)$, this latter difference is at most $\frac{v(o)}{w_{i}}$ for some object $o\in A_{j}$. Therefore, the difference is at most $\frac{V}{w_{i}}$. ∎
Theorem 4.7.
For additive identical valuations, Algorithm 2 computes a WEF-able allocation $A$ in $O(mn+m\log(m))$ time, where $\forall i\in N$: $s_{i}\leq V$. Therefore, the total subsidy required is at most $(n-1)V$.

### v4

@@ 3.2 Computing a WEF-able allocation
Assume $P_{i}(A)=(i\ldots j)$ is the highest-cost path from $i$ in $G_{A,w}$. Note that $\ell_{i}(A)=cost_{A}(P_{i}(A))$.
@@
$cost_{A}(C)=\ell_{i}(A)+cost_{A}(j,i).$
By Section 3.1, $cost_{A}(C)\leq 0$, thus,
$\ell_{i}(A)\leq-cost_{A}(j,i)\leq z.$
Therefore, $s_{i}=w_{i}\ell_{i}(A)\leq w_{i}z$. ∎
We use Section 3.2 with a modified valuation function.
We prove that an allocation that is WEF-able for the original valuations is also WEF-able for the modified valuations (Section 3.2), and that the maximum subsidy required by each agent for the original valuations is bounded by the subsidy required for the modified valuations (Section 3.2).
Next, we demonstrate that under the modified valuations, the cost of each edge is at least $-V$. Finally, by Section 3.2, we conclude that the maximum subsidy required for any agent $i\in N$ is $w_{i}V$ for the modified valuation $\bar{v}$ (Section 3.2) and for the original valuations $v$ as well.
@@
By Section 3.1, it is sufficient to prove that all directed cycles in the envy graph (with the modified valuations and unit weights) have non-positive total cost. We prove a stronger claim: in every round $t$, the total modified-cost added to every directed cycle $C$ is non-positive. Let $A^{t}$ be the allocation computed by Algorithm 1 at iteration $t$. Suppose, contrary to our assumption, that there exists a cycle $C=(i_{1},...,i_{r})$ and a round $t$ in which the modified-cost added to $C$ is positive. To simplify notation, we consider $i_{1}$ as $i_{r+1}$. This implies that
@@
Case 1: All arcs $i_{j}\to i_{j+1}$ in $C$ have
$\bar{v}_{i_{j}}(A_{i_{j+1}}^{t})=\frac{v_{i_{j}}(A_{i_{j+1}}^{t})}{w_{i_{j+1}}}$
(in particular, this holds for $t=T$). In this case, inequality (1) implies
$\overline{cost}_{A^{t}}(C)=\sum_{j=1}^{r}\frac{v_{i_{j}}(A_{i_{j}}^{t})}{w_{i_{j+1}}}-\frac{v_{i_{j}}(A_{i_{j}}^{t})}{w_{i_{j}}}>0.$
Combined with Section 3.1, this contradicts Section 3.2, which states that $A^{t}$ is WEF-able.
Case 2: All arcs $i_{j}\to i_{j+1}$ in $C$ have
$\bar{v}_{i_{j}}(A_{i_{j+1}}^{t})=\frac{v_{i_{j}}(A_{i_{j}}^{t+1})}{w_{i_{j}}}.$
In this case, inequality (1) implies
$\sum_{j=1}^{r}\frac{v_{i_{j}}(A_{i_{j}}^{t+1})}{w_{i_{j}}}>\sum_{j=1}^{r}\frac{v_{i_{j}}(A_{i_{j}}^{t})}{w_{i_{j}}}.$
Notice that all the items in $A_{j_{1}}^{t+1},...,A_{j_{r}}^{t+1}$ are available at iteration $t$, which contradicts the optimality of $\{A_{i}^{t}\}$.
Case 3: Some arcs $i_{j}\to i_{j+1}$ in $C$ satisfy Case 1 and the other arcs satisfy Case 2. Let $l\geq 1$ be the number of arcs in $C$ that satisfy Case 2. We decompose $C$ into a sequence of $l$ edge-disjoint paths, denoted $P_{1},...,P_{l}$, such that the last node of each path is the first node of the next path, and in each path, only the last edge satisfies Case 2. Formally, suppose that some path contains $k\geq 1$ agents, denoted as $i_{1},...,i_{k}$, and $k-1$ arcs. Then for each $21\leq j\leq k-$ 2, $\bar{v}_{i_{j}}\left(A_{i_{j+1}}^{t}\right)=\frac{v_{i_{j}}\left(A_{i_{j+1}}^{t}\right)}{w_{i_{j+1}}}$ and $\bar{v}_{i_{k-1}}\left(A_{i_{k}}^{t}\right)=\frac{v_{i_{k-1}}\left(A_{i_{k-1}}^{t+1}\right)}{w_{i_{k-1}}}$. Since $\overline{cost}_{A^{t}}(C)>0$, there exits a path $P=(i_{1},...,i_{k})$ where $\overline{cost}_{A^{t}}(P)>0$, which implies that:
$0<\sum_{j=1}^{k-1}\left(\bar{v}_{i_{j}}(A_{i_{j+1}}^{t})-\bar{v}_{i_{j}}(A_{i_{j}}^{t})\right)=\sum_{j=1}^{k-2}\left(\frac{v_{i_{j}}(A_{i_{j+1}}^{t})}{w_{i_{j+1}}}-\frac{v_{i_{j}}(A_{i_{j}}^{t})}{w_{i_{j}}}\right)+\frac{v_{i_{k-1}}(A_{i_{k-1}}^{t+1})}{w_{i_{k-1}}}-\frac{v_{i_{k-1}}(A_{i_{k-1}}^{t})}{w_{i_{k-1}}}$.
The rest of the proof is similar to the proof of Section 3.2. We construct another allocation $B^{t}$ randomly as follows:
@@
$\mathbb{E}\left[\sum_{i\in N}\left(v_{i}(B_{i}^{t})-v_{i}(A^{t}_{i})\right)\right]=\sum_{1\leq j\leq k-2}\left(\mathbb{E}\left[v_{i_{j}}(o_{i_{j+1}}^{t})\right]-\mathbb{E}\left[v_{i_{j}}(o_{i_{j}}^{t})\right]\right)+$
$\mathbb{E}\left[v_{i_{k-1}}(o_{i_{k-1}}^{t+1})\right]-\mathbb{E}\left[v_{i_{k-1}}(o_{i_{k-1}}^{t})\right]=\sum_{1\leq j\leq k-2}\frac{v_{i_{j}}(A^{t}_{i_{j+1}})}{w_{i_{j+1}}}-\frac{v_{i_{j}}(A^{t}_{i_{j}})}{w_{i_{j}}}+$
$\frac{v_{i_{k-1}}(A_{i_{k-1}}^{t+1})}{w_{i_{k-1}}}-\frac{v_{i_{k-1}}(A_{i_{k-1}}^{t})}{w_{i_{k-1}}}.$
@@
Given Section 3.2, for each $i,j\in N$,
$\bar{v}_{i}(A_{j})-\bar{v}_{i}(A_{i})\geq\frac{v_{i}(A_{j})}{w_{j}}-\frac{v_{i}(A_{i})}{w_{i}}.$
@@
By Section 3.2, the allocation $A$ is WEF-able under the valuations $\bar{v}$. Together with Section 3.2, if for each $i,j\in N$ it holds that $\bar{v}_{i}(A_{j})-\bar{v}_{i}(A_{i})\geq-V$, the subsidy required for agent $i\in N$ is at most $w_{i}V$ for $\bar{v}$.
@@
$\sum_{t\in[T-1]}\max\Big{\{}\frac{v_{i}(A_{j}^{t})}{w_{j}},\frac{v_{i}(A_{i}^{t+1})}{w_{i}}\Big{\}}+\frac{v_{i}(A_{j}^{T})}{w_{j}}-\sum_{t\in[T]}\frac{v_{i}(A_{i}^{t})}{w_{i}}\geq$
@@
We are now prepared to prove the main theorem. {theorem} For additive valuations and integer entitlements, Algorithm 1 computes in polynomial time a WEF-able allocation, where the subsidy to each agent is at most $w_{i}V$ and the total subsidy is at most $(W-w_{1})V$.
@@
By Section 3.2, $A$ is WEF-able under the original valuations. Combined with Section 3.2 and Section 3.2 , $A$ is also WEF-able under the modified valuations and requires a subsidy of at most $w_{i}V$ for each agent $i\in N$.
Section 3.2, implies that under the original valuations, the required subsidy for each agent $i\in N$ is at most $w_{i}V$. By Section 3.1, there is at least one agent who requires no subsidy, so the required total subsidy is at most $(W-w_{1})V$.
For the runtime analysis, the most computationally intensive step in Algorithm 1 is solving the maximum integral flow of minimum cost in $G^{\prime}$. The flow network $G^{\prime}$ consists of at most $n+m+2$ nodes and at most $n+m+mn$ arcs. By goldberg1989finding, this can be done in time polynomial in $n,m$:
$O\left(\left(n+m+2\right)\left(n+m+mn\right)\log\left(n+m+2\right)\right.\\ \left.\min\{\log\left(\left(n+m+2\right)V\right),\left(n+m+mn\right)\log\left(n+m+2\right)\}\right).$
The WEF condition is invariant to multiplying the weight vector by a scalar. This can be used in two ways:
(1) If the weights are not integers, but their ratios are integers, we can still use Algorithm 1. For example, if $w_{1}=1/3$ and $w_{2}=2/3$ (or even if $w_{i}$’s are irrational numbers such as $w_{1}=\sqrt{2}$ and $w_{2}=2\sqrt{2}$), Algorithm 1 works correctly by resetting $w_{1}=1$ and $w_{2}=2$.
(2) If the weights are integers with greatest common divisor (gcd) larger than 1, we can divide all weights by the gcd to get a better subsidy bound:
{lemma}
For additive valuations and integer entitlements, there exists an algorithm that computes in polynomial time a WEF-able allocation where the subsidy to each agent is at most $w_{i}V/\operatorname{gcd}(\mathbf{w})$ and the total subsidy is at most $(W-w_{1})V/\operatorname{gcd}(\mathbf{w})$, where $\operatorname{gcd}(\mathbf{w})$ is the greatest common divisor of all the $w_{i}$.
Proof.
Algorithm 1 works correctly, even if we divide each $w_{i}$ by the greatest common divisor of $w_{i}$’s. In other words, letting $d={\rm gcd}(w_{1},...,w_{n})$, $w^{\prime}_{i}=w_{i}/d$, $W^{\prime}=W/d$, and running Algorithm 1 with $w^{\prime}_{i}$’s, we get the bound $(W^{\prime}-w^{\prime}_{\min})V$ of the total subsidy. ∎
@@ 4 WEF Solutions for Additive Identical Valuations
In this section This section deals with the case where all agents have identical valuations, that is, $v_{i}\equiv v$ for all $i\in N$.
@@
The value per unit entitlement of each agent $i<n$ is $V(w_{i}-1)/w_{i}$, and for agent $n$ it is $V$. Therefore, to avoid envy, each agent $i<n$ should receive a subsidy of $w_{i}\left(1-\frac{w_{i}-1}{w_{i}}\right)V=V$ and the total subsidy required is $(n-1)V.$
@@
[…]
Given an instance with identical valuations, let $A$ be any allocation, and denote by $P$ any path in the weighted envy-graph of $A$ between agents $i,j\in N$. Then,
$cost_{A}(P)=\frac{v(A_{j})}{w_{j}}-\frac{v(A_{i})}{w_{i}}.$
This is because the path cost is
$\sum_{(h,k)\in P}cost_{A}(h,k)=\sum_{(h,k)\in P}\frac{v(A_{k})}{w_{k}}-\frac{v(A_{h})}{w_{h}},$
and the latter sum is a telescopic sum that reduces to the difference of its last and first element.
Using Section 4, it is easy to prove the following Lemma. {lemma} With identical valuations, every allocation is WEF-able.
@@
$\begin{split}&cost_{A}(C)=\\ &\sum_{k=1}^{r-1}\left(\frac{v(A_{i_{k+1}})}{w_{i_{k+1}}}-\frac{v(A_{i_{k}})}{w_{i_{k}}}\right)+\frac{v(A_{i_{1}})}{w_{i_{1}}}-\frac{v(A_{i_{r}})}{w_{i_{r}}}=\\ &\sum_{k=1}^{r}\left(\frac{v(A_{i_{k}}}{w_{i_{k}}}-\frac{v(A_{i_{k}}}{w_{i_{k}}}\right)=0.\end{split}$ (2)
Hence, by condition (b) of Section 3.1, $A$ is WEF-able. ∎
@@
The algorithm traverses the items in an arbitrary order. At each iteration it selects the agent that minimizes the expression $\frac{v(A_{i}\cup{\{o\}})}{w_{i}}$, with ties broken in favor of the agent with the larger $w_{i}$, and allocates the next item to that agent. Intuitively, this selection minimizes the likelihood that weighted envy is generated.
1:Instance $(N,M,v,\mathbf{w})$ with additive identical valuations.
2:WEF-able allocation $A$ with total required subsidy of at most $(n-1)V$.
3:$A_{i}\leftarrow\emptyset,\forall i\in N$
4:for $o:1$ to $m$ do
5:     $I^{t}\leftarrow\arg\min_{i\in N}\frac{v\left(A_{i}\cup\{o\}\right)}{w_{i}}$
6:     $i^{t}\leftarrow\max_{i\in I^{t}}\left(i\right)$
7:     Add $o$ to $A_{i^{t}}$
8:end for
9:return $A$
The following example illustrates Algorithm 2:
[…]
@@
    In order to mitigate envy, $s_{1}=\frac{6}{7}$ and $s_{2}=0$.
Example 4 illustrates that the resulting allocation may not beWEF(1,0)Algorithm 2 might allocate all items to the agent with the highest entitlement. However, the outcome is always WEF$(0,1)$:
{proposition}
@@
We prove by induction that at each iteration, $A$, the resulting allocation from Algorithm 2, satisfies WEF$(0,1)$. The claim is straightforward for the first iteration. Assume the claim holds for the $(t-1)$-th iteration, and prove it for the $t$-th iteration. Let $o$ be the item assigned in this iteration and $i^{t}$ be the agent receiving this item. Agent $i^{t}$ satisfies WEF$(0,1)$ due to the induction hypothesis. For $j\neq i^{t}$, by the selection rule, $\frac{v(A_{i^{t}})}{w_{i^{t}}}\leq\frac{v(A_{j}\cup\{o\})}{w_{j}}$. This is exactly the definition of WEF$(0,1)$. ∎
Based on Section 4 and Section 4, we show that each path starting at agent $i\in N$ under $WEF(0,1)$ allocation is bounded by $\frac{V}{w_{i}}$. {proposition} With identical additive valuations, for every WEF$(0,1)$ allocation $A$, $\ell_{i}(A)\leq\frac{V}{w_{i}}$, for all $i\in N$.
@@
For each agent $i\in N$, denote the highest-cost path starting at $i$ in that graph by $P_{i}(A)=(i,...,j)$ for some agent $j\in N$. Then by Section 4, $\ell_{i}(A)=cost_{A}(P_{i}(A))=\frac{v(A_{j})}{w_{j}}-\frac{v(A_{i})}{w_{i}}$.
From the definition of WEF$(0,1)$, Algorithm 2 implies that this difference is at most $\frac{v(o)}{w_{i}}$ for some object $o\in A_{j}$. Therefore, the difference is at most $\frac{V}{w_{i}}$. ∎
{theorem}
For additive identical valuations, Algorithm 2 computes a WEF-able allocation $A$ in $O(mn)$ time, where $\forall i\in N$: $s_{i}\leq V$. Therefore, the total subsidy required is at most $(n-1)V$.